OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08. A. 1.5 B. 2 C. 2.4 D. 3

OpenStudy (amistre64):

t = (A/P - 1)/r t = (168/150 - 1)/.08

OpenStudy (anonymous):

@gageallen15 this formula is not simple interest!

OpenStudy (anonymous):

I = p * r * t is the right one!

OpenStudy (anonymous):

The other is Compound Interest: A = P ( 1 + r /n ) ^(nt)

OpenStudy (anonymous):

@gageallen15 where are you?